For each of the following, compute the present value: (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Present Value Years Interest Rate Future Value

$ 10 6 % $ 17,928

2 11% 42,117

14 14% 796,382

19 13% 653,816
Answer #1
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars Amount
a) PV = 17928/((1.06)^10)               10,010.90
b) PV = 42117/((1.11)^2)               34,183.10
c) PV = 796382/((1.14)^14)             127,190.16
d) PV = 653816/((1.13)^19)               64,115.81